find vertex, focus, and equation of directrix of y^2 = 36x
Answer by lwsshak3(11628)

find vertex, focus, and equation of directrix of y^2 = 36x
This is an equation of a parabola that opens rightwards.
Its standard form of equation: (y-k)=4p(x-h), (h,k)=(x,y) coordinates of the vertex.
For given equation: y^2=36x
vertex:(0,0)
axis of symmetry: y=0 or x-axis
4p=36
p=9
focus: (9,0) (p-distance to the right of the vertex on the axis of symmetry)
directrix: x=9 (p-distance to the left of the vertex on the axis of symmetry)